# LLM as GNN: Graph Vocabulary Learning for Text-Attributed Graph Foundation Models

### v2

@@ Appendix E Implementation Details
We provide additional details to facilitate reproducibility. In the graph understanding module, the number of layers in the prompt-based GNN is selected from ${1,2,3,4}$, and the first-order neighbor sampling ratio from ${0.3,0.6,1.0}$. In the graph inference module, we fine-tune the LLM with a learning rate of 3e-4 and a batch size of 4. To mitigate potential biases introduced by task-specific prompts, we designed a prompt pool for each task requirement and randomly selected prompts during instruction construction to enhance robustness. We employed a standard early-stopping strategy during training: if the performance metric on the validation set did not improve over a fixed number of consecutive epochs (determined based on the dataset), we halted training to prevent overfitting. For other hyperparameters of the compared methods, we referred to the original papers and carefully tuned them to suit each dataset.
Appendix F Prompt Length Limitation Discussion
@@
In dense TAGs, high-degree nodes pose a challenge for prompt-based GNNs due to potentially excessive prompt lengths, which can exceed the input limits of current LLMs and impair model efficiency. To address this issue, we analyze in the following aspects that balances efficiency, scalability, and representation quality.
F.1 Neighbor Sampling in Scalable GNNs
@@
Following standard GNN practices, we sample a fixed number of neighbors per node to reduce prompt length and improve robustness. As shown in Table 11, varying the sampling ratio across three datasets reveals that selecting 60% of neighbors yields the best performance, confirming its effectiveness. Meanwhile, limiting the maximum number of sampled neighbors is a common practice in scalable GNNs. For example, GraphSAGE samples 25 and 10 neighbors at the first and second hops, respectively, while PinSAGE selects from $\{10,20,50\}$. Inspired by these practices, we choose from $\{10,20,30,40\}$ and empirically set 20 as the cap, ensuring prompts remain within LLM input limits while preserving sufficient structural context.
F.2 Token Capacity of LLMs
@@
Our implementation leverages GPT-4o mini, which supports up to 128k input tokens per call. This capacity allows a single prompt to encode information from over 300 neighbors, given that each neighbor’s textual summaries occupies approximately 400 tokens on average (for the first round of message propagation). As shown in Table 12, only a small fraction of nodes (<0.25%) in a few datasets exceed this limit. For these rare cases, we have applied neighbor sampling and truncation in Section F.1, effectively bounding prompt length without degrading representation quality.
F.3 Prompt Design
In our prompt-based GNN framework, we design prompts that guide the LLM to aggregate information from a node’s sampled neighbors and produce a concise yet expressive representation. This process is inherently iterative: with each layer, the prompt becomes more compressed as semantic and structural information is distilled into fewer tokens. Empirically, the average prompt length decreases from approximately 400 tokens in the first layer, to  125 tokens in the second, and eventually to  10 tokens per node on the Cora dataset.
Importantly, each node is processed independently, and its prompt length depends solely on its local neighborhood, i.e., its degree, rather than on global graph size. This localized design ensures scalability, as prompt lengths remain bounded even in large graphs, allowing our approach to maintain efficiency without sacrificing representational capacity.
@@ F.3 Prompt Design
Table 11: The performance with varying neighbor sampling ratios acorss three datasets.
@@
Table 12: Analysis of High-Degree Nodes and Average Connectivity
Metric Cora Citeseer PubMed Ogbn-arxiv History Photo WikiCS
Avg. Degree 4.01 2.84 13.77 36.90 18.07 12.11 4.50
High-degree (>300) Nodes 0 0 0 0.20% 0.25% 0.19% 0
Appendix G Permutation Sensitivity
We evaluate the sensitivity of our prompt-based GNNs (using Flan-T5) to input order by evaluating two variants: (a) shuffle nodes, which randomizes the order of neighbors, and (b) shuffle tokens, which injects position-specific tokens to nodes and shuffle them instead. As illustrated in Table 13, regardless of the shuffle method, both variants result in only minor fluctuations in node classification accuracy, indicating that prompt-based GNNs with the subsequent fine-tuning are robust to input permutation and maintain stable performance.
Table 13: Permutation sensitivity of prompt-based GNNs.
      Variant       Cora       Citeseer
      shuffle nodes       90.64 $\pm$1.12       83.79 $\pm$ 0.89
      shuffle tokens       90.55 $\pm$1.03       84.28 $\pm$1.06
Appendix H Case Study
H.1 Textual Representations in Prompt-based GNNs
In this part, we select two representative cases in citation networks and demonstrate their layer-by-layer refinement of our prompt-based GNN. Specifically, we provide the textual representations at each layer, including Round 0 as initial features. More importantly, we also collect its one-hop nodes and highlighted the source and relevant information below. From our empirical studies, we have the following observations. Overall, it is evident that the verbose textual representations are progressively refined to concise textual presentations. Meanwhile, the core semantics become increasingly clear throughout the process, until a short sequence composed of several natural language tokens at the last round. Furthermore, we notice that we effectively incorporate the key ideas of some neighboring nodes, as reflected in the refined textual representations after each aggregation-update operation. More specifically, the neighboring information is naturally fused to the central node within the text space, boosting their connections and improving the expressiveness of the central node. This study further illustrates the explicit message passing through our proposed prompt-based GNN in the text space, which captures the graph semantics and structures simultaneously.
H.1.1 Case A
@@
Round 1: 1st Layer Representations. This bootstrap method improves neural network evaluations in NYSE forecasting, comparable to traditional error estimation methods. It clears and simplifies data structures, enhancing accuracy and predicting probability distribution.
/***"comparable to traditional error estimation methods" comes from the paper: A comparison of some error estimates for neural network models. ***/
/***"It clears and simplifies data structures" comes from the paper: The Observer-Observation Dilemma in Neuro-Forecasting: Reliable Models From Unreliable Data Through CLEARNING. ***/
Round 2: 2nd Layer Representations. This bootstrap-inspired method refines neural network evaluations for NYSE forecasting, highlighting local error bars from nonlinear regression analysis, improving accuracy with data simplification techniques.
/***"local error bars from nonlinear regression analysis" comes from the paper: Learning Local Error Bars for Nonlinear Regression. ***/
@@
H.1.2 Case B
Raw Text Attributes. Paper Title: Insulin dependent diabetes in Asians. Abstract: Type 1 diabetes is said to be extremely rare in children in India, where diabetes treated with insulin may be due to chronic pancreatic disease or malnutrition. To see whether typical type 1 diabetes occurred in Asian children in the United Kingdom, all known Asian children with diabetes in industrial West Yorkshire were ascertained. A total of 17 such children were studied; of these, seven were from three multiplex families and two fathers from these families had diabetes. All children were ketosis prone and developed diabetes while resident in the UK. There were significant increases in HLA-B8 and HLA-DR3 and increases in HLA-DR4 and HLA-DR3/DR4, while HLA-B15 was absent. Islet cell antibodies, either IgG or complement fixing, were present in four of 18 subjects tested, all of whom had disease of short duration. The prevalence of type 1 diabetes in Asian children aged 15 years or less in West Yorkshire was 36/100,000, assuming complete ascertainment. It is concluded that typical type 1 diabetes may occur in Asian children and this condition may be more common in families who have migrated to the UK.
Round 0: Textual Feature Initialization. The study examined type 1 diabetes prevalence in 17 Asian children in West Yorkshire, UK, focusing on family history. Seven were from multiplex families, and all were ketosis-prone. Significant increases in HLA-B8, HLA-DR3, and HLA-DR4 were found, with no HLA-B15. Islet cell antibodies were present in four of 18 subjects. The diabetes prevalence was 36/100,000, suggesting it may be more common in migrant families.
Round 1: 1st Layer Representations. The study examined type 1 diabetes in Asian children in West Yorkshire, UK. Environmental factors and migration may raise incidence, unlike the low incidence of insulin-dependent diabetes in Karachi. Increases in HLA-B8, HLA-DR3, and HLA-DR4 were found.
/***"Environmental factors and migration may raise incidence" comes from the paper: Evidence for an environmental effect in the aetiology of insulin dependent diabetes in a transmigratory population. ***/
/***"low incidence of insulin-dependent diabetes in Karachi" comes from the paper: Incidence of insulin dependent diabetes mellitus in Karachi, Pakistan. ***/
Round 2: 2nd Layer Representations. The study examined type 1 diabetes in Asian children in West Yorkshire, UK. Migration may raise incidence, unlike the low incidence in Karachi. Increases in HLA markers were found, similar to North and South Indian diabetics.
/***"Migration may raise incidence" comes from the paper: HLA-DR antigen frequencies in a North Indian type I diabetic population. ***/
/***"North and South Indian diabetics" comes from two papers: HLA-DR antigen frequencies in a North Indian type I diabetic population and HLA, complement C2, C4, properdin factor B and glyoxalase types in South Indian diabetics. ***/
Round 3: 3rd Layer Representations. Type 1, diabetes. Migration and HLA markers linked to increased diabetes incidence.
H.2 Language-based IDs vs. Key Words
To provide further insights, we leverage external information to validate the superiority of PromptGFM in capturing the core semantics of nodes. Specifically, we extract key words of papers from citation datasets (Cora, Citeseer, Obgn-arxiv, and PubMed) and compare them with their language-based IDs in our universal graph vocabulary.
Table 14 summarizes the key words and language-based IDs of selected papers, along with their titles and URLs for reference. Overall, it is evident that there are strong semantic relevance between the language-based IDs and keywords. For example, regarding the paper titled Distributed Protocols at the Rescue for Trustworthy Online Voting, the key words have appeared within its language-based ID, suggesting that PromptGFM has effectively captured its core semantics through our prompt-based GNN. In addition, in Committees providing EJR can be computed efficiently, where the title is less indicative of the content, the language-based ID still aligns perfectly with the corresponding key words, such as efficient computation and rules. This finding demonstrates that PromptGFM not only effectively captures the core idea without relying on the title, but also filters relevant semantics from neighboring nodes to enhance its own representations. Overall, our language-based IDs accurately capture and extend the semantics of the nodes, making them well-suited to form a universal graph vocabulary.
Table 14: The comparison between language-based IDs generated by prompt-based GNNs as part of the graph vocabulary and keywords extracted from the original papers in citation datasets. To prevent data leakage, we exclude these keywords from model training. The observed similarity confirms that our prompt-based GNN effectively captures the essential semantics of nodes. Furthermore, our language-based IDs encode richer information than keywords by preserving high-order signals.
Paper Language-based ID Key Words URL
Modular Verification of Interrupt-Driven Software Modular verification of interrupt-driven software using abstract interpretation Software, Abstract Interpretation, Feasibility Verification arXiv:1709.10078
Parsimonious Data: How a single Facebook like predicts voting behaviour in multiparty systems Predicting voting behavior using Facebook likes in multiparty systems Facebook Likes, Voter Intention, Machine Learning, Multiparty System arXiv:1704.01143
A Fast Noniterative Algorithm for Compressive Sensing Using Binary Measurement Matrices Fast noniterative algorithm for compressive sensing with binary matrices Compressive Sensing, Deterministic Methods arXiv:1708.03608
Optimization of Battery Energy Storage to Improve Power System Oscillation Damping Battery storage optimization improves power system oscillation damping Battery Energy Storage System, Oscillation Damping arXiv:1811.10213
Neural Variational Hybrid Collaborative Filtering Neural Variational Hybrid Collaborative Filtering improves recommendation performance Collaborative Filtering, VAE, Recommendation System arXiv:1810.05376v6
Interpretable Neural Networks for Predicting Mortality Risk using Multi-modal Electronic Health Records Predicting mortality risk using interpretable multi-modal neural network Mortality Risk Prediction, Clinical Data arXiv:1901.08125
A New Approach to Distributed Hypothesis Testing and Non-Bayesian Learning: Improved Learning Rate and Byzantine-Resilience Distributed hypothesis testing with Byzantine resilience using Bayesian update Bayesian Learning, Byzantine Resilience arXiv:1907.03588
Accurate and Efficient Hyperbolic Tangent Activation Function on FPGA using the DCT Interpolation Filter Efficient hyperbolic tangent activation function using DCTIF Hyperbolic Tangent, Activation Function arXiv:1609.07750
Distributed Protocols at the Rescue for Trustworthy Online Voting Trustworthy online voting with distributed blockchain protocols Distributed Voting, Distributed Protocols arXiv:1705.04480
Committees providing EJR can be computed efficiently Efficient computation of approval-based multi-winner voting rules Approval-Based Voting, Multi-Winner Elections arXiv:1704.00356
Creatism: A deep-learning photographer capable of creating professional work Creatism: deep learning system for artistic photography creation Creatism, Evaluation of Photographic Quality, Deep Learning arXiv:1707.03491
Relation of familial patterns of coronary heart disease, stroke, and diabetes to subclinical atherosclerosis: the multi-ethnic study of atherosclerosis Family history beyond early-onset heart disease impacts atherosclerosis Family History, Coronary Heart Disease, Stroke doi.org/10.1097/GIM.0b0
13e31818e639b
Glycemic index, glycemic load, and risk of type 2 diabetes Benefits of low-GI diet in type 2 diabetes Diabetes, Prevention doi.org/10.1093/ajcn/76/1.
274S
Decreased insulin responsiveness of glucose uptake in cultured human skeletal muscle cells from insulin-resistant nondiabetic relatives of type 2 diabetic families Inherited defects contribute to insulin resistance in diabetes Insulin Resistance, Inherited Factors doi.org/10.2337/diabetes.
49.7.1169
Quantitative histopathological studies of the extramural coronary arteries from Type 2 (non-insulin-dependent) diabetic patients Histopathological study of coronary arteries in diabetic patients Histopathology, Diabetes Mellitus doi.org/10.1007/BF0027
4798
Metabolic control and diet in Finnish diabetic adolescents Factors influencing metabolic control in diabetic adolescents Diabetes Mellitus, Adolescent doi.org/10.1111/j.1651-2227.1992.tb12212.x
H.3 Impact Statement
We highlight the following concrete cases to better demonstrate the broad applicability of our approach.
$\bullet$ In recommender systems, different scenarios involve distinct graph structures and scenario-specific semantics. Instead of training scenario-specific models, our foundation model could act as a general-purpose recommendation backbone that unifies multiple scenarios in text space via a co-training paradigm. This enables LLMs to capture rich and transferable semantic relationships across scenarios (e.g., user preferences, item affinities, contextual behaviors). Specifically, (a) our foundation model alleviates cross-dataset inconsistencies and facilitates rapid adaptation to new scenarios and tasks. (b) it is particularly effective in cold-start issues and long-tail recommendation challenges, where existing methods often struggle due to rigid architectures and heavy dependence on labeled data.
$\bullet$ In biomedical knowledge graphs, entities (e.g., genes, proteins, drugs, diseases) and their complex relations (e.g., inhibition, expression, association) form semantically rich graphs. Our approach achieves cross-resource knowledge integration (e.g., DrugBank, CTD, DisGeNET) and supports inductive reasoning across biomedical subdomains, enabling fast adaptation to new tasks and unseen relation types with minimal supervision. This is particularly valuable in biomedical settings, where knowledge resources are often fragmented, relation annotations are sparse, and entity semantics vary significantly across datasets. In this context, our approach is well suited for low-resource scenarios such as rare disease research, novel drug discovery, and cross-database knowledge transfer.

### v3

@@ A.3. Implementation Details
We provide additional details to facilitate reproducibility. In the graph understanding module, the number of layers in the prompt-based GNN is selected from $\{1,2,3,4\}$, and the first-order neighbor sampling ratio from $\{0.3,0.6,1.0\}$. In the graph inference module, we fine-tune the LLM with a learning rate of 3e-4 and a batch size of 4. To mitigate potential biases introduced by task-specific prompts, we designed a prompt pool for each task requirement and randomly selected prompts during instruction construction to enhance robustness. We employed a standard early-stopping strategy during training: if the performance metric on the validation set did not improve over a fixed number of consecutive epochs (determined based on the dataset), we halted training to prevent overfitting. For other hyperparameters of the compared methods, we referred to the original papers and carefully tuned them to suit each dataset.
Figure 7. Graph understanding module via prompt-based GNNs. We prompt LLMs to achieve fine-grained reproduction of traditional GNN workflow, refining verbose textual representations into concise yet meaningful ones. In the prompt, neighbor sampling (see Equation 3) is in blue, the aggregation-update mechanism (see Equation 4) in blue, and the optimization in red.
Appendix B Prompt Design
This section provides templates of prompts, including prompt-based GNNs and task-oriented prompts.
Prompt for node initialization. The title of the paper is ¡the title of the paper¿, the abstract of the paper is ¡the abstract of the paper¿. Please summarize the paper. /*** This prompt varies depending on the dataset. The present instance is for citation datasets only. ***/
Prompt for each layer of GNN replication. Given the central node ¡l-th round textual representation of the central node¿. The selected one-hop neighbors are [¡ l-th round of node #1¿, ¡l-th round node #2¿, … ,¡l-th round node #N¿]. Please aggregate neighbor nodes and update a concise yet meaningful representation for the central node. Note connected nodes should share similar semantics and vice versa.
Prompt for node classification. ¡central node language-based ID¿ has 1-hop connections with […, ¡1-hop neighbor language-based IDs¿, …], and it also has 2-hop connections with […, ¡2-hop neighbor language-based IDs¿, …]. Which category should ¡central node language-based ID¿ be classified as?
Prompt for discriminative link prediction. ¡central node language-based ID ¿ has 1-hop connections with […, ¡1-hop neighbor language-based IDs¿, …], and it also has 2-hop connections with […, ¡2-hop neighbor language-based IDs¿, …]. Among ¡central node language-based ID¿ and ¡negative sampling node language-based ID¿, which node will be connected to ¡central node language-based ID¿ ?
Prompt for generative link prediction. ¡central node language-based ID¿ has 1-hop connections with […, ¡1-hop neighbor language-based IDs¿, …], and it also has 2-hop connections with […, ¡2-hop neighbor language-based IDs¿, …]. Which node should be connected to ¡central node language-based ID¿?
Table 14. Analysis of High-Degree Nodes and Average Connectivity
Metric Cora Citeseer PubMed Ogbn-arxiv History Photo WikiCS
Avg. Degree 4.01 2.84 13.77 36.90 18.07 12.11 4.50
High-degree (¿300) Nodes 0 0 0 0.20% 0.25% 0.19% 0
Appendix C Further Exploration
C.1. Prompt Length Limitation Discussion
@@
In dense TAGs, high-degree nodes present a challenge for prompt-based GNNs, as their long prompts may exceed LLM input limits and hinder efficiency. To address this issue, we examine strategies that balance efficiency, scalability, and representational fidelity.
C.1.1. Neighbor Sampling in Scalable GNNs.
@@
Following standard practices, we adopt a neighbor sampling strategy in each GNN layer to reduce prompt length and enhance robustness. As shown in Table 15, sampling 60% of neighbors yields the best overall performance across datasets. Meanwhile, limiting the maximum sampled neighbors is a common practice in scalable GNNs. For example, GraphSAGE samples 25 and 10 neighbors at the first and second hops, while PinSAGE selects from $\{10,20,50\}$. Following this rationale, we explore $\{10,20,30,40\}$ and empirically set 20 as the cap, balancing LLM input constraints with sufficient structural coverage.
C.1.2. Token Capacity of LLMs.
@@
Our implementation leverages GPT-4o mini, which supports up to 128k input tokens per call. This capacity allows a single prompt to encode information from over 300 neighbors, given that each neighbor contributes about 400 tokens in the first propagation round. As shown in Table 14, only a small fraction of nodes (¡0.25%) in a few datasets surpass this capacity. For these rare high-degree nodes, our neighbor sampling and truncation mechanism can bound prompt length, ensuring broad scalability to dense graphs without information loss.
C.1.3. Prompt Design.
Prompt-based GNNs guide the LLM to aggregate and update neighbor information through structured prompts, producing increasingly compact yet expressive representations. Across layers, the average length shrinks from  400 to  10 tokens as semantic and structural information are progressively distilled. Since each node is processed independently and depends only on its local degree, prompt size remains bounded regardless of graph scale, ensuring both efficiency and scalability to large graphs.
@@
Table 15. The performance with varying neighbor sampling ratios acorss three datasets.
@@
Table 16. The effectiveness of adaptive prompting that alleviates inter-domain negative transfer.
Source Pre-training Co-training
Acc ($\uparrow$) M-F1 ($\uparrow$) Acc ($\uparrow$) M-F1 ($\uparrow$)
History Photo w/o w/ w/o w/ w/o w/ w/o w/
39.12 40.23 39.84 40.13 90.67 90.82 91.82 91.93
46.24 47.44 47.41 47.98 88.52 90.27 89.39 90.22
44.21 45.11 44.93 46.02 85.34 88.13 85.57 88.47
45.18 47.90 45.65 47.26 86.22 89.02 87.03 89.16
C.2. Negative Transfer Discussion
C.2.1. Why does performance drop?
Cross-domain negative transfer mainly stems from domain shifts, including structure heterogeneity, semantic shift in text, and label space mismatch (especially for node classification), rather than hard vocabulary conflicts. In contrast, PromptGFM adopts a unified language-based vocabulary and pure language interface, which enables natural cross-domain transferability and avoids token collision in principle.
C.2.2. Adaptive Prompting Exploration.
[…]
C.3. Case Study
We visualize the layer-by-layer refinement of node representations and identify which neighbor nodes contribute to each stage. Textual representations evolve from verbose descriptions into compact, high-density token sequences (8-10 tokens). After the first layer, nodes absorb neighbor semantics (e.g., ”error estimation”); after the second round, they integrate broader two-hop context (e.g., ”nonlinear regression analysis”). These observations demonstrate the explicit message passing behavior of the prompt-based GNN in text space, effectively capturing both graph semantics and structural dependencies.
@@ C.3. Case Study
Round 1: 1st Layer Representations. This bootstrap method improves neural network evaluations in NYSE forecasting, comparable to traditional error estimation methods [A]. It clears and simplifies data structures [B], enhancing accuracy and predicting probability distribution.
/***[A] comes from the paper: A comparison of some error estimates for neural network models. ***/
/***[B] comes from the paper: The Observer-Observation Dilemma in Neuro-Forecasting: Reliable Models From Unreliable Data Through Clearning. ***/
Round 2: 2nd Layer Representations. This bootstrap-inspired method refines neural network evaluations for NYSE forecasting, highlighting local error bars from nonlinear regression analysis [C], improving accuracy with data simplification techniques.
/***[C] comes from the paper: Learning Local Error Bars for Nonlinear Regression. ***/